OpenStudy (anonymous):

can anyone help? show that of all the rectangles with a given area, the one with the smallest perimeter is a sqaure

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

are u allowed to use calculus?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok.... so let P=2x+2y, where the rectangle has length x, and width y...

OpenStudy (anonymous):

so we need to minimize P....

OpenStudy (anonymous):

area of the rectangle is given, so A=xy . this is our secondary equation that we'll need to substitute into the primary equation (perimeter equation)

OpenStudy (anonymous):

solving for y, we have y=A/x

OpenStudy (anonymous):

substituting into the perimeter equation we have: P = 2x + 2(A/x) \(\large P = 2x + 2Ax^{-1} \) now you differentiate P....

OpenStudy (anonymous):

if i continue from where u are i have p'=a

OpenStudy (anonymous):

no... you did not differentiate correctly...

OpenStudy (anonymous):

P'=\[\pm \sqrt{A}\]

OpenStudy (anonymous):

sorry X not P'

OpenStudy (anonymous):

\(\large P=2x+2Ax^{-1} \) \(\large P'=2-2Ax^{-2} \)

OpenStudy (anonymous):

now set that derivative equal to zero and solve for x....

OpenStudy (anonymous):

alright....i see it now

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\(\large P'=2-2Ax^{-2} \) \(\large 0=2-2Ax^{-2} \) \(\large 0=1-Ax^{-2} \) \(\large Ax^{-2}=1 \) \(\large x^{-2}=\frac{1}{A} \) \(\large x^{2}=\frac{A}{1} =A\) \(\large x=\sqrt{A} \) since x = squareroot A, y must be squareroot A also... in other words, x=y....

OpenStudy (anonymous):

since x=y, the rectangle is a square...

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw... :)